drop down includes:

towards the top of the page

into the page

out of the page

towards the bottom of the page
Answer #1

at point C

magnetic force = q * v * B * sin(theta)

magnetic force = 1.602 *10^-19 * 0.34 * 5.8 *10^5 * sin((90+ 15.2) degree)

magnetic force = 8.282 *10^-15 N

the magnetic force is 3.049 *10-14 N

----------------------------

Using right hand rule

the direction of force is OUT of the page